2. An investment company is planning upgrades to its website. The company believes 36% of its customers access the website through a smartphone. a) If the company takes a random sample of 200 recent visitors to the website, and checks the IP address to see if access was through a smartphone, give the properties (and provide a sketch) for the sampling ample proporti smartphone b) Find the probability the sample will yield less than 30% having accessed the website from a smartphone.
Answer #1

2)a) n = 200

       p = 0.36

\mu_{\widehat p} = p = 0.36

\sigma_{\widehat p} = sqrt(p(1 - p)/n)

      = sqrt(0.36 * (1 - 0.36)/200)

     = 0.0339

b) P(\widehat p < 0.3)

= P((\widehat p - \mu_{\widehat p} )/\sigma_{\widehat p} < (0.3 - \mu_{\widehat p} )/\sigma_{\widehat p})

= P(Z < (0.3 - 0.36)/0.0339)

= P(Z < -1.77)

= 0.0384